is 3 let the X N UNIFCol), and dowrity function of y =-1n x. Then what y where non zero!
Answer #1

The probability density function (PDF) of X is

f(x) = 1,0<x<1

The cumulative distribution function (CDF) of X is

F(x) = 1,0 <r<1

Now,

The cumulative distribution function (CDF) of Y = -ln X is

Fyly) = P(Y SY) = P(-In X <y) = P(X>ey)=1- P(X <ey)

→ Fy(y) = 1- P(X <ey)

Fr(y) = 1- Fxley)

Fy(y) = 1-e,y> 0

The probability density function (PDF) of Y = -ln X is

(9) = (1) = ਆ --91 = 1,9 > 0